Which of the following has highest percentage of oxygen by mass? (1) h20 (2) co2 (3)caco3 (4) c2h5oh

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
Answer:

caco3 has the highest percentage of oxygen by mass

Thanks
Krisha
2 years ago
CaCO3 has molecular weight 100g.
100g of CaCO3 contain (16*3) 80g of oxygen.
% by mass of oxygen= 80/100*100 =80% 
In H2O, % by mass of oxygen=16/18*100
=88% 
In CO2 ,% by mass of oxygen=32/44*100
=72%
In C2H5OH, % by mass of oxygen=16/46*100 =34%
H2O has higher % of oxygen by mass.